A calorie is a unit of heat or energy and it equals about 4.2 J where 1 J = 1 kgm2 s-2. Suppose we employ a system of units in which the unit of mass equals a kg, the unit of length equals j8 m, the. unit of time is ys. Show that a calorie has a magnitude 4.2  α-1 β-2 γ2 in terms of the new units.


1
Expert's answer
2021-02-15T05:43:47-0500

"1kg_{new}=\alpha*1{kg}"

"1{kg}=\frac{1}{\alpha}*1kg_{new}"

"1m_{new}=\beta*1m"

"1m=\frac{1}{\beta}*1m_{new}"

"1m^2=\frac{1}{\beta^2}*1m^2_{new}"

"1s_{new}=\gamma*1s"

"1s=\frac{1}{\gamma}*1s_{new}"

"1s^{-2}={\gamma}^2*1s_{new}"

"4.2J=4.2[kgm^2s^{-2}]=4.2*1kg*1m^2*1s^{-2}="

"4.2*\frac{1}{\alpha}*1kg_{new}*\frac{1}{\beta^2}*1m^2_{new}*{\gamma}^2*1s_{new}="

"4.2*{\alpha}^{-1}*{\beta^{-2}}*{\gamma}^2*1kg_{new}*1m^2_{new}*1s_{new}="

"4.2{\alpha}^{-1}{\beta^{-2}}{\gamma}^2[kg_{new}m^2_{new}s_{new}]"

"\text{Answer:}\ 4.2{\alpha}^{-1}{\beta^{-2}}{\gamma}^2[kg_{new}m^2_{new}s_{new}]"